OpenStudy (anonymous):

I will be your fan and give a medal 1. -3x - 3y = 6 -2x + 5y = -24 Find the solution to the system of equations shown above by graphing. a) x = 2, y = -4 b) x = 3, y = -4 c) x = 2, y = -3 d) x = -4, y = 2

OpenStudy (anonymous):

What you do is turn the equations into y-intercept form and then graph them. The solutions will be where the x intercept is and the y intercept is.

OpenStudy (anonymous):

So basically...

OpenStudy (anonymous):

Equation 1: -3x - 3y = 6 -3x - 3y = 6 +3x +3x <--- Add 3x to both sides, canceling out one and adding . the other shown below. -3y = 6 + 3x /-3 /-3 <---Divide -3 (to leave y by itself) from both sides. Make . sure you divide it like (6 + 3x)/-3 y = x - 2 ^Now graph that equation on the graph.

OpenStudy (the_fizicx99):

You don't have to graph the equation @HaleyRoadcap You can simply plug in the for y into the other equation and solve for x. Plotting works too, but you would need both to be in y = mx + b form ans see where they would intersect.

OpenStudy (anonymous):

And do the same with the other equation

OpenStudy (anonymous):

@tHe_FiZiCx99 is right, as well. I just always found it easier to do the longer way.

OpenStudy (anonymous):

ok, thx

OpenStudy (the_fizicx99):

That's the graphical way, I prefer it algebraically.